the following set of lead determinations on seperate aliquots of a pooled blood serum were reported 103,106,107 and 114 meq/L.one value appears suspect.should the suspect be retained or discarded at 95% confidence level



1
Expert's answer
2022-05-23T10:55:03-0400

X1= 103 X2= 106 X3= 107 X4= 114 N=4

We use Dixon's Q-test to solve this problem:

1. The values comprising the set of observations under examination are arranged in ascending order:

X1 < X2 < X3 < X4

2. The statistic experimental Q-value (Qexp) is calculated. This is a ratio defined as the difference of thesuspect valuefrom itsnearestone divided by the range of the values:

Qexp = (X2 - X1) / (X4 - X1) => (106-103) / (114-103) = 3/8 = 0.375

3. The obtained Qexp value is compared to a critical Q-value (Qcrit) found in tables. This critical value should correspond to the confidence level (CL) we have decided to run the test (CL=95%).

The following figure shows a table of critical values ​​of Q.




4. If Qexp > Qcrit, then the suspect value can be characterized as an outlier and it can be rejected, if not, the suspect value must be retained and used in all subsequent calculations.

Qexp = 0,375

Qcrit = 0,829

Qexp < Qcrit hence, the suspect should be detained

ANSWER: suspect should be detained